OpenStudy (anonymous):

A speeder passes a parked police car at a constant speed of 36.7m/s. At that instant, the police car starts from rest with a uniform acceleration of 2.83m/s. How much time passes before the speeder is overtaken by the police car? how far does the speeder travel before being overtaken by the police car in meters

OpenStudy (anonymous):

So the speeder has position x = 36.7t the police car has position x = (1/2)gt^2 (where g is the acceleration of the police car). when the police car is overtaken, their x positions are the same so we can equate them 36.7t = x = (1/2)gt^2 now solve this for t by dividing t from each side 36.7 = (1/2)gt (2 x 36.7)/g = t t = 25.9 seconds to find the distance they will be, plug this t back into either equation. ill use the first x = 36.7t = 36.7 x 25.9 = 950m